If x^2 = 6x-9 , then what value is 5x^2+x-8 ? The hint is to solve for x first.
Answer by sofiyac(983)

If x^2 = 6x-9 , then what value is 5x^2+x-8 ? The hint is to solve for x first.
5(6x-9)+x-8=5x^2+x-8
5(6x-9)=5x^2
6x-9=x^2
x^2-6x+9=0
(x-3)(x-3)=0
so x=3
5x^2+x-8=5*9+3-8=45-5=40 The answer is 40